Accelerator Update for December 1, 2008 - December 3, 2008
Monday Dec 1
The day shift began with Operations monitoring the Tevatron (TeV) store 6601 (46.42E30) and the Recycler stash (173.9E10), with the Antiproton Source stacking (35.5E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the detectors for the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments in standby due to a cold box controls problem, and with the MiniBooNE experiment taking beam. NuMI remained down to replace their horn.
At 8:22 AM, the Cryo Coordinator reported that he needed to turn off all the TeV cold compressors in order to maintain the helium inventory. Operations terminated store 6601. Experts reported later that they had resolved the PLC issues that caused the cold box problems.
Operations prepared CDF and D0 for access at 9:54 AM.
CDF came out of access at 11:12 AM.
At 11:49 AM, Operations reported noise problems with a Linac quadrupole power supply (QPS109) to the Linac experts.
Linac experts completed their work on QPS109 and Operations resumed sending beam to Pbar and MiniBooNE at 2:02 PM.
Booster experts reported at 2:50 PM that the problem with RF station #18 (BRF18) was probably due to its power amplifier in the tunnel. An access would be needed to make repairs.
Cryo system experts reported at 2:55 PM that all the TeV cold compressors were running.
Operations reported at 3:53 PM that the Recycler had stashed antiprotons eight times during the day shift.
Operations put the TeV through a dry squeeze at 4:21 PM.
Operations put the TeV through a wet squeeze at 5:43 PM.
Operations put the TeV into shot setup at 6:09 PM.
Operations established store 6605 at 7:45 PM with an initial luminosity of 339.19E30. This store had the second highest luminosity.
Operations reported at midnight that the Recycler had stashed five times during the evening shift.
Tuesday Dec 2
The midnight shift began with Operations monitoring store 6605 (164.43E30) and the stash (141.6E10), with Pbar stacking (44.2E10), and with MiniBooNE taking beam.
At 5:40 AM, Operations reported more problems with the noisy Linac quadrupole (QPS109).
Operations reported at 7:55 AM that the Recycler had stashed eight times during the midnight shift.
The Neutron Therapy Facility reported at 9:56 AM that they had competed treatment on their last patient for the day.
At 10:21 AM, Linac experts began repairs on L:QPS109. Operations put Pbar into standby mode and halted the MI ramp for the work.
Operations resumed sending beam to Pbar and MiniBooNE at 10:47 AM.
Operations terminated store 6605 at 11:12 AM.
Operations put the TeV into shot setup at 11:23 AM.
Operations reported at 1:20 PM that LCW experts were replacing the pressure transducer for the MI-20 LCW pump #2. They were also making up water to the MI LCW system.
Operations established store 6606 at 1:35 PM with an initial luminosity of 335.6E30. This store had the third highest luminosity: 1. 11/26/08 – store 6591 344.3E30, 2. 12/1/08 – store 6601 339.19E30.
A Pbar expert reported at 3:15 PM that he would be conducting an antiproton study during the next Recycler stash and then a Pbar Debuncher stochastic cooling study that would inhibit stacking. The studies ended at 4 PM due to problems with a MI power supply (Q713).
Operations reported at 4 PM that the Recycler had stashed three times during the day shift.
At 4:32 PM, MIRF14 tripped off for the first of two times in 30 minutes. It reset. A RF expert investigated.
Operations reported at midnight that the Recycler had stashed seven times during the evening shift.
Wednesday Dec 3
The midnight shift began with Operations monitoring store 6606 (99.78E30) and the stash (224.4E10), with Pbar stacking (31.5E10), and with MiniBooNE taking beam.
Operations terminated store 6606 at 7:02 AM.
Operations prepared CDF for an access at 7:18 AM.
At 7:40 AM, Operations reported that the stacking rate had dropped due to a problem with the Pbar PLC. An expert investigated.
A Pbar expert reported at 7:54 AM that the PLC power supply had failed. They switched to a backup power supply.
Operations reported at 8:02 AM that the Recycler had stashed eight times during the midnight shift.
The Plans for Wednesday
The plans for today are to allow CDF to finish their access and then go into shot setup.
Comments and Suggestions
|last modified 01/05/2009 email Fermilab|
|Security, Privacy, Legal|